# A Survey on Recent Advances in LLM-Based Multi-turn Dialogue Systems

### v1

@@ 1.1. What is Multi-turn Dialogue System?
Multi-turn dialogue systems that can generate natural and meaningful responses to communicate with humans are a long-term goal of artificial intelligence (AI). Such human-computer interaction task has attracted increasing attention from academia and industry due to their potential impact and attractive commercial value. The multi-turn dialogue task can be regarded as a sequence-to-sequence task, which generates the system response $\mathcal{S}=(s_{1},s_{2},...s_{t})$ from the user message $\mathcal{U}=(u_{1},u_{2},...u_{t})$, where $u_{t}$ and $s_{t}$ are the user message and system response in the $t$-th round, respectively.
@@ 1.2. Why a Survey on LLM-based Multi-turn Dialogue System?
Arora et al. (arora2013dialogue, ) provided an overview of dialogue systems and introduced various dialogue system frameworks. However, this survey treated the dialogue system as a generic system rather than categorizing it into TOD and ODD systems and did not encompass deep learning models. Chen et al. (chen2017survey, ) categorized dialogue systems into TOD and ODD systems, discussing the application of deep learning techniques in both types of dialogue systems. Nevertheless, this survey did not dive into multi-turn dialogue systems based on pre-trained LLMs. The review written by Ni et al. (ni2023recent, ) covered multi-turn dialogue systems based on pre-trained LLMs, but this study did not provide detailed insights into LLMs and methods for adapting them to downstream sub-tasks. In contrast, Qin et al. (qin2023end, ) provided a more comprehensive exploration of the application of pre-trained LLMs in specific-target dialogue scenarios. However, the focus of this paper primarily centered on end-to-end task-oriented multi-turn dialogue systems.
Our paper aims to provide a state-of-the-art overview of LLM-based multi-turn dialogue systems, before which we will provide a comprehensive exposition of existing pre-trained LLMs and methodologies employed to adapt these models for downstream tasks. This investigation is anticipated to appeal to broad audiences in both academia and industry, encompassing researchers and practitioners alike.
@@ 1.3. Contribution of this Survey
In this paper, we provide a comprehensive review of methods, evaluation metrics and datasets for LLM-based multi-turn dialogue. The contribution of our paper can be summarized as follows:
@@
(2) To provide a detailed exposition on state-of-the-art multi-turn dialogue datasets and evaluation metrics.
@@
The rest of this survey is organized as follows. In Sec.2, we provide a detailed exposition of prevalent LLMs. From Sec.3 to Sec.4, we will give a comprehensive introduction to methods for adapting LLMs to downstream tasks. In Sec.5, we present important methods of TOD, including pipeline-based methods and end-to-end methods. State-of-the-art methods of ODD are proposed in Sec.6. In Sec.7 and Sec.8, we introduce some relevant datasets and evaluation metrics. Besides, some problems and challenges of LLM based multi-turn dialogue are proposed in Sec.9. Finally, we conclude our survey in Sec.10.
@@ 2. GENERAL METHODS
The vanilla Transformer architecture (vaswani2017attention, ), a sequence-to-sequence model, has emerged as a foundational framework for diverse LLMs, utilizing encoders and decoders with self-attention mechanisms as its core components, thanks to its exceptional parallelism and capacity. Based on the masking methods utilized by various attention mechanisms in the model, the current LLMs can be divided into three categories, i.e., Encoder-Decoder, Decoder-only, and Encoder-only. The decoder-only category further includes distinctions such as causal decoders and prefix decoders, illustrated in Figure 1.
@@
Figure 1. The matrix comparison of attention mask patterns between decoder-only and encoder-decoder architectures. The matrix uses dark cells to allow for self-attention of input elements $j$ at the output time step $i$, while light cells restrict this attention. The left panel represents the full input attention, the middle panel refers to preventing future input reliance, and the right panel combines causal masking with a prefix for partial input sequence fully-visible masking. (raffel2020exploring, )
@@ 2.1. Decoder-only Transformer Architecture
The decoder-only model (raffel2020exploring, ), functioning independently without an encoder, can act as a language model designed primarily for next-step prediction (liu2018generating, ; radford2018improving, ; al2019character, ). During the training of language models, decoder-only model takes on the task of generating the target sequence.
@@ 2.1.1. Causal Decoder
The causal decoder architecture employs unidirectional attention masking to ensure that each input token can only attend to past tokens and itself. Both input and output tokens are processed in a similar manner within the decoder. The schematic for this architecture is depicted in the middle panel of Figure 1.
@@
The Generative Pre-trained Transformer (GPT) model has garnered significant attention and interest in the field of NLP (ye2023comprehensive, ), and the technical evolution of the GPT series models is illustrated in Figure 2. Positioned as a cutting-edge technology built upon the Transformer architecture, the versatility and robust performance of the GPT model establish it as a universal solution for various NLP tasks.
Figure 2. A brief illustration for the technical evolution of GPT-series models. We created this flowchart primarily relying on information from research papers, blog articles, and official APIs provided by OpenAI. The solid line represents explicit evidence of the evolutionary path between the two models (e.g., the official statement that a new model is developed based on a base model), while the dashed line denotes a relatively weaker evolutionary relationship.
GPT-1.
Proposed by the OpenAI team, GPT-1 (radford2018improving, ), short for Generative Pre-trained Transformer 1 (vaswani2017attention, ), serves as the foundational model for the GPT-series, establishing the key architecture and the fundamental principle of modeling natural language text, specifically predicting the next word. GPT-1 adopts a decoder-only Transformer architecture, implementing a semi-supervised approach for language understanding through a combination of unsupervised pre-training and supervised fine-tuning.
GPT-2.
GPT-2 (radford2019language, ), is an extension of the GPT-1 architecture (radford2018improving, ), where the parameter scale is increased to 1.5 billion. Trained on the extensive WebText dataset, GPT-2 aims to perform diverse tasks through unsupervised language modeling, eliminating the need for explicit fine-tuning with labeled data. GPT-2 utilizes a probabilistic form for multitask solving, denoted as $p(output|input,task)$, predicting the output based on input and task information. This approach is reminiscent of similar methods found in (mccann2018natural, ). Leveraging a multi-layer self-attentive mechanism, GPT-2 achieves fully-connected cross-attention to the entire context in a computationally efficient manner.
GPT-3.
GPT-3 (brown2020language, ) employs attention mechanisms, allowing the model to selectively focus on segments of input text it deems most relevant and adopts an autoregressive approach, leveraging the transformative power of the Transformer architecture.
GPT-3.5.
The GPT-3.5 models, also known as InstructGPT (ouyang2022training, ), are developed based on the code-davinci-002 underscoring the efficacy of training on code data to improve the reasoning capacity of GPT models (zhao2023survey, ). InstructGPT, is a language model fine-tuned on GPT-3 (brown2020language, ) using a combination of supervised learning and reinforcement learning based on human feedback (RLHF). GPT-3.5 is utilized to develop the chatbot product ChatGPT (openai_chatgpt, ; lock2022ai, ). The gpt-3.5-turbo is the most capable and cost-effective GPT-3.5 model. GPT-3.5 shows improvements in maintaining truthfulness and reducing toxic outputs, demonstrating that fine-tuning with human feedback is an effective approach to enhance the alignment of language models with human intent.
GPT-4.
GPT-4 (openai2023gpt4, ), is a multimodal LLM, accepting both image and text inputs and produces text outputs. Despite being less capable than humans in various real-world scenarios, GPT-4 showcases human-level performance across a range of professional and academic benchmarks. Notably, GPT-4 has improved calibration compared to GPT-3.5 (ouyang2022training, ), exhibiting enhanced accuracy in predicting the correctness of its answers. GPT-4’s versatility extends to ChatGPT (openai_chatgpt, ), where it can process images as inputs.
ChatGPT.
ChatGPT (openai_chatgpt, ), short for Chat Generative Pre-trained Transformer, represents a chatbot innovation. Leveraging a robust LLM, it empowers users to shape and guide conversations according to specific criteria such as length, format, style, level of detail, and language (lock2022chatgpt, ). ChatGPT is built upon either GPT-3.5 (ouyang2022training, ) or GPT-4 (openai2023gpt4, ), and is fine-tuned for conversational applications using a combination of supervised learning and reinforcement learning (gertner2023wikipedia, ). ChatGPT Plus (chatgpt-openai, )is a version of ChatGPT backed by GPT-4. Users of ChatGPT Plus have the ability to upload images, and mobile app users can engage in conversations with the chatbot (roose2023chatgpt, ).
GPTs.
GPTs (openaigpts, ) represent an innovative feature enabling users to customize versions of ChatGPT (openai_chatgpt, ) for specific purposes. These personalized GPTs assist users in enhancing daily efficiency, optimizing performance in specific tasks, and simplifying work or home activities. Users can share their creations with others to promote the utility of these purpose-built GPTs. GPTs prioritize privacy and safety by ensuring user control over data within the ChatGPT environment. Interactions with GPTs remain confidential, with no sharing of user conversations with the creators.
@@
LLaMA (touvron2023llama, ), short for Large Language Model Meta AI, is a family of LLMs, released by Meta AI (llama2023, ). Due to its openness and effectiveness, LLaMA attracts significant attention from the research community, and many efforts are devoted to fine-tuning or continually pre-training its different model versions for implementing new models or tools. The technical evolution of the LLaMA series models is illustrated in Figure 3.
Figure 3. A brief illustration for LLaMA series.
LLAMA.
LLaMA (touvron2023llama, ), a collection of foundation language models ranging from 7B to 65B parameters, is trained on trillions of tokens, and show that it is possible to train state-of-the-art models using publicly available datasets exclusively, without resorting to proprietary and inaccessible datasets. In particular, LLaMA-13B outperforms GPT-3 (brown2020language, ) (175B) on most benchmarks, and LLaMA-65B is competitive with the best models at that time, Chinchilla-70B (hoffmann2022training, ) and PaLM-540B (rae2021scaling, ).
LLAMA2.
LLaMA2 (llama2announcement, ), the next iteration in the LLaMA series, has been unveiled through a collaborative effort between Meta and Microsoft. Llama 2 is a collection of pre-trained and fine-tuned LLMs ranging in scale from 7 billion to 70 billion parameters. The Llama 2 pre-trained models, trained on 2 trillion tokens, feature double the context length compared to Llama. The release of LLaMA2 includes model weights and starting code for pre-trained and fine-tuned LLaMA language models (Llama Chat, Code Llama).
LLAMA2 CHAT.
LLaMA2 includes both foundational models and models fine-tuned for dialog, called LLaMA 2-Chat (touvron2023llama2, ). Llama 2-Chat, are optimized for dialogue use cases. The models outperform open-source chat models on most benchmarks tested, and, based on human evaluations for helpfulness and safety, may be a suitable substitute for closed-source models.
CODE LLAMA.
Code Llama (rozière2023code, ) contains three versions with different parameter amounts, i.e., 700 million parameter version, 1.3 billion parameter version and 34 billion parameter version. When training the basic model, first initialize the weights with the Llama 2 model (llama2announcement, ) with the same number of parameters, and then train on the code dataset of 500 billion words. Meta has also fine-tuned the trained basic model in two different styles: Python expert version (plus 100 billion additional words), and instruction fine-tuning version which can understand natural language instructions.
@@ 2.1.2. Prefix Decoder
The prefix decoder structure (raffel2020exploring, ) modifies the causal encoder’s masking mechanism to enable bidirectional attention on prefix tokens while maintaining unidirectional attention on generated tokens. Similar to the encoder-decoder paradigm, this allows bidirectional encoding of prefix sequences and auto-regressive generation of output tokens, sharing the same parameters during both encoding and decoding phases. Prominent models utilizing the prefix decoder architecture include U-PaLM, GLM-130B, and others in large-scale prefix encoders. The schematic for this architecture is depicted in the right panel of Figure 1.
GLM
GLM (du2021glm, ), short for General Language Model, is a comprehensive pre-training framework. The technical evolution of the GLM series models is illustrated in Figure 4. The architecture modifications in GLM include rearranging the order of layer normalization and the residual connection, using a single linear layer for output token prediction, and replacing ReLU activation functions with GeLUs. GLM unifies pre-training objectives for various tasks under the common framework of autoregressive blank infilling, employing mixed attention masks and novel 2D position encodings.
Figure 4. A brief illustration for GLM series.
ChatGLM series
ChatGLM (zeng2022glm, ) is a bilingual model incorporating question-answering, multi-turn dialogue, and code generation. Built upon GLM-130B (zeng2022glm, ), ChatGLM follows the design principles of ChatGPT (openai_chatgpt, ).
ChatGLM-6B.
ChatGLM-6B (zeng2022glm, ), as the inaugural ChatGLM conversational model, builds upon the training insights from GLM-130B. It resolves issues in the implementation of 2D RoPE positional encoding and adopts a conventional FeedForward Network(FFN) structure.
ChatGLM2-6B.
ChatGLM2-6B (thudm2022chatglm2-6b, ), incorporates a blend of target functions within the GLM framework, undergoing pre-training with 1.4 trillion Chinese and English identifiers and aligning with human preferences. Leveraging FlashAttention technology, ChatGLM2-6B extends the base model’s context length from 2K (ChatGLM-6B) to 32K, employing an 8K context length during dialogue training. With the integration of Multi-Query Attention technology, ChatGLM2-6B achieves more efficient inference speed and lower GPU memory usage, boasting a 42% improvement in inference speed compared to its predecessor in the official model implementation.
ChatGLM3-6B.
ChatGLM3-6B (THUDM2023chatglm3, ) introduces a newly designed Prompt format, accommodating normal multi-turn conversations as well as native support for complex scenarios such as tool invocation (Function Call), code execution (Code Interpreter), and Agent tasks.
@@ 2.2. Encoder-only Transformer Architecture
Unlike decoder-only and encoder-decoder LLMs that utilize autoregressive regression, the encoder-only LLMs emphasize comprehension of input content and generation of task-specific outputs.
@@ 2.2.1. BERT Series
BERT (devlin2018bert, ), short for Bidirectional Encoder Representations from Transformers, is a language model based on the transformer architecture, known for its significant improvement over previous state-of-the-art models. The brief illustration for BERT series is shown in Figure 5.
Figure 5. A brief illustration for BERT series.
BERT
BERT (devlin2018bert, ) namely a bidirectional Transformer-based encoder, is a LLM introduced by the Google AI Language team (bertopensourcing, ). It incorporates masked language modeling, allowing pre-training to capture interactions between left and right context words. Recent advancements, such as extended training duration, parameter tying across layers, and span masking instead of individual words, have demonstrated improved performance. Notably, BERT’s auto-regressive predictions limit its effectiveness for generation tasks.
RoBERTa
RoBERTa (liu2019roberta, ), namely Robustly optimized BERT (devlin2018bert, ) approach, improves upon BERT (devlin2018bert, ) through straightforward modifications, such as longer training with larger batches, removal of the next sentence prediction objective, and dynamic changes to the masking pattern, and employs a more extensive byte-level Byte Pair Encoding (BPE) vocabulary. However, unlike BERT, RoBERTa streamlines its training process by omitting the Next Sentence Prediction (NSP) task and focusing solely on optimizing the Masked Language Model (MLM) task. This approach enhances the model’s ability to learn bidirectional contextual information in language.

### v2

@@ 1.1. What is Multi-turn Dialogue System?
The task of dialogue generation can be broadly classified into single-turn and multi-turn dialogues. Single-turn dialogue systems generate responses based solely on the current user input, without considering previous interactions. In contrast, multi-turn dialogue systems need to maintain context over several turns, incorporating both the current user message and the dialogue history to generate relevant, coherent, and contextually appropriate responses. This introduces several challenges, such as handling long-range dependencies, maintaining consistency across turns, and managing ambiguous or incomplete information. These challenges make multi-turn dialogue systems significantly more complex and demanding than single-turn systems.
Multi-turn dialogue systems aim to generate natural and meaningful responses that facilitate effective communication with humans. Achieving this goal has been a long-term focus in the field of artificial intelligence (AI), as such systems hold significant potential for advancing human-computer interaction. The growing interest in multi-turn dialogues, both in academia and industry, is driven by their promising applications and considerable commercial value.
In the context of multi-turn dialogue, the task can be viewed as a sequence-to-sequence problem, where the system generates a response sequence $\mathcal{S}=(s_{1},s_{2},...s_{t})$ based on the input sequence $\mathcal{U}=(u_{1},u_{2},...u_{t})$, with each $u_{t}$ and $s_{t}$ representing the user’s message and the system’s response at the $t$-th turn, respectively. This task requires not only understanding the current user input but also maintaining and utilizing the dialogue history to generate coherent and contextually relevant responses.
@@ 1.2. Why a Survey on LLM-based Multi-turn Dialogue System?
Traditional multi-turn dialogue systems face several critical challenges that have long hindered their practical applications. First, these systems often struggle with maintaining contextual coherence across multiple turns, frequently failing to track long-term dependencies and maintain consistent dialogue states. Second, they typically require extensive domain-specific knowledge engineering and carefully curated training data, making them costly to develop and challenging to scale across domains. Third, these systems often exhibit limited flexibility in handling unexpected user inputs and complex dialogue scenarios that deviate from their training patterns. Moreover, they frequently struggle with understanding implicit context, managing diverse linguistic variations, and generating contextually appropriate responses that maintain both semantic accuracy and natural fluency across extended conversations.
LLM-based approaches offer promising solutions to these longstanding challenges (zhao2023survey, ). Through their massive-scale pre-training on diverse corpora, LLMs demonstrate superior capabilities in understanding and maintaining dialogue context across multiple turns (touvron2023llama2, ). Their strong generalization ability significantly reduces the dependency on extensive task-specific training data, enabling more efficient cross-domain deployment (gao2023chat, ). Furthermore, LLMs exhibit remarkable flexibility in handling diverse dialogue scenarios and understanding implicit context, thanks to their sophisticated semantic understanding and generation capabilities. The capacity for few-shot learning in LLMs opens up new possibilities for rapid adaptation to novel domains and tasks without extensive retraining (madotto2020language, ).
Arora et al. (arora2013dialogue, ) provided an overview of dialogue systems and introduced various dialogue system frameworks. However, this survey treated the dialogue system as a generic system rather than categorizing it into TOD and ODD systems and did not encompass deep learning models. Chen et al. (chen2017survey, ) categorized dialogue systems into TOD and ODD systems, discussing the application of deep learning techniques in both types of dialogue systems. Nevertheless, this survey did not dive into multi-turn dialogue systems based on pre-trained LLMs. The review written by Ni et al. (ni2023recent, ) covered multi-turn dialogue systems based on pre-trained LLMs, but this study did not provide detailed insights into how LLMs fundamentally transform the architecture and capabilities of dialogue systems. In contrast, Qin et al. (qin2023end, ) provided a more comprehensive exploration of the application of pre-trained LLMs in specific-target dialogue scenarios. However, the focus of this paper primarily centered on end-to-end task-oriented multi-turn dialogue systems, leaving gaps in understanding the broader implications of LLMs across different dialogue paradigms.
Our survey aims to provide a comprehensive survey of methods, evaluation approaches, and datasets for LLM-based multi-turn dialogue, with an emphasis on both task-oriented and open-domain dialogue systems. We systematically review the methodologies for adapting LLMs to downstream dialogue tasks, analyze the current state and limitations of existing approaches, and discuss potential future research directions in this field.
@@ 1.3. Contribution of this Survey
Figure 1. The overall diagram of this article
As shown in Figure 1, we produced a diagram for this article to help readers familiarize the overall structure. In this paper, we provide a comprehensive review of methods, evaluation metrics and datasets for LLM-based multi-turn dialogue. The contribution of our paper can be summarized as follows:
@@
(2) To systematically analyze how LLM-based approaches address traditional challenges in multi-turn dialogue systems and transform the landscape of human-machine interaction;
(3) To provide a detailed exposition on state-of-the-art multi-turn dialogue datasets and evaluation metrics;
@@
The rest of this survey is organized as follows. In Sec.2, we provide a detailed exposition of prevalent LLMs and methods for adapting LLMs to downstream tasks. In Sec.3, we present important methods of TOD, including pipeline-based methods and end-to-end methods. State-of-the-art methods of ODD are proposed in Sec.4. In Sec.5 and Sec.6, we introduce some relevant datasets and evaluation metrics. Besides, some problems and challenges of LLM based multi-turn dialogue are proposed in Sec.7. Finally, we conclude our survey in Sec.8.
@@ 2. GENERAL METHODS
The vanilla Transformer architecture (vaswani2017attention, ), a sequence-to-sequence model, has emerged as a foundational framework for diverse LLMs, utilizing encoders and decoders with self-attention mechanisms as its core components, thanks to its exceptional parallelism and capacity. Based on the masking methods utilized by various attention mechanisms in the model, the current LLMs can be divided into three categories, i.e., Encoder-Decoder, Decoder-only, and Encoder-only. The decoder-only category further includes distinctions such as causal decoders and prefix decoders.
@@ 2.1.1. Causal Decoder
The causal decoder utilizes unidirectional attention masking, ensuring each token attends only to past tokens and itself. Both input and output tokens are processed similarly.
@@
The Generative Pre-trained Transformer (GPT) models, based on the Transformer architecture, have become a cornerstone in NLP, demonstrating impressive versatility and performance across various tasks. GPT-1 (radford2018improving, ), the first in the series, established a semi-supervised approach that combined unsupervised pre-training with supervised fine-tuning, using a decoder-only Transformer for next-word prediction. Building on this, GPT-2 (radford2019language, ) expanded the model to 1.5 billion parameters, utilizing unsupervised learning for diverse tasks, with a multi-layer self-attention mechanism that efficiently captures cross-context dependencies. GPT-3 (brown2020language, ) further refined the model by enhancing its autoregressive approach, leveraging advanced attention mechanisms to improve performance on a wide range of tasks. The next iteration, GPT-3.5 (ouyang2022training, ) (InstructGPT), incorporated reinforcement learning from human feedback (RLHF), improving accuracy, reducing harmful outputs, and enabling the development of ChatGPT (openai_chatgpt, ), a conversational model. GPT-3.5 is available in the cost-effective GPT-3.5-turbo version, which powers ChatGPT, making it more efficient and user-friendly. GPT-4 (openai2023gpt4, ), the multimodal model, supports both text and image inputs, excelling in accuracy and achieving human-level performance across professional and academic benchmarks. It powers ChatGPT Plus, which allows users to upload images and access the model on mobile devices.
GPTs (openaigpts, ), an innovative feature of the ChatGPT platform, enable users to customize versions of the model for specific tasks while maintaining control over their data and ensuring privacy. These personalized GPTs can be shared with others to enhance productivity and optimize performance in various domains.
@@
The LLaMA (touvron2023llama, ) models, released by Meta, are a family of open-source language models that have gained significant attention in the research community for their effectiveness and adaptability. LLaMA models, ranging from 7B to 65B parameters, are trained on publicly available datasets, showcasing that state-of-the-art models can be built without relying on proprietary datasets. LLaMA-13B has outperformed GPT-3 (brown2020language, ) on most benchmarks, and LLaMA-65B has been shown to be competitive with other leading models, such as Chinchilla-70B (hoffmann2022training, ) and PaLM-540B (rae2021scaling, ). The next generation, LLaMA2 (llama2announcement, ), includes models ranging from 7B to 70B parameters, trained on 2 trillion tokens, and offers double the context length compared to LLaMA, making it more suitable for complex tasks. LLaMA2 includes fine-tuned models like LLaMA Chat and Code LLaMA, further enhancing its utility. LLaMA2-Chat (touvron2023llama2, ), optimized for dialogue applications, outperforms other open-source chat models in benchmarks and can serve as a viable alternative to closed-source models. Additionally, Code LLaMA (rozière2023code, ), a version fine-tuned for programming tasks, has been trained on a massive code dataset and is available in multiple sizes, with models fine-tuned for specific languages and instruction-following capabilities.
@@ 2.1.2. Prefix Decoder
The prefix decoder structure (raffel2020exploring, ) modifies the causal encoder’s masking mechanism, enabling bidirectional attention for prefix tokens while maintaining unidirectional attention for generated tokens. This allows shared parameters for both encoding and decoding phases, as seen in models like U-PaLM, GLM-130B, and other large-scale prefix encoders.
GLM (du2021glm, ), a pre-training framework, introduces several architectural modifications such as rearranged layer normalization, a single linear layer for token prediction, and GeLU activations. It unifies multiple tasks under a common autoregressive infilling objective, using mixed attention masks and 2D position encodings. Building on this, ChatGLM (zeng2022glm, ), derived from GLM-130B (zeng2022glm, ), is designed for bilingual tasks like question-answering, dialogue, and code generation, following the principles of ChatGPT (openai_chatgpt, ). The first conversational version, ChatGLM-6B (zeng2022glm, ), resolves issues with 2D RoPE positional encoding and employs a standard FeedForward Network (FFN) structure.
The subsequent model, ChatGLM2-6B (thudm2022chatglm2-6b, ), improves upon its predecessor by incorporating 1.4 trillion Chinese and English identifiers for pre-training and extending the context length to 32K tokens. It also integrates FlashAttention and Multi-Query Attention technologies, achieving a 42% boost in inference speed and reducing GPU memory usage. Finally, ChatGLM3-6B (THUDM2023chatglm3, ) introduces a new prompt format, supporting not only multi-turn conversations but also complex tasks like tool invocation, code execution, and agent-based operations.
@@ 2.2. Encoder-only Transformer Architecture
Unlike decoder-only and encoder-decoder LLMs that rely on autoregressive regression, encoder-only LLMs like BERT focus on understanding input content and generating task-specific outputs.
@@ 2.2.1. BERT Series
BERT (devlin2018bert, ) (Bidirectional Encoder Representations from Transformers) is a transformer-based language model that significantly improves previous state-of-the-art models. It uses masked language modeling to capture bidirectional context, but its autoregressive prediction limits its effectiveness for generation tasks. BERT’s performance has been further enhanced by techniques like extended training duration, parameter tying across layers, and span masking. RoBERTa (liu2019roberta, ) builds on BERT by removing the next sentence prediction objective and using longer training with larger batches, dynamic masking, and a more extensive byte-level Byte Pair Encoding (BPE) vocabulary, optimizing the model for better bidirectional contextual learning.